How many moles of ammonium sulfate contain 6.0×1025 atoms of N

N?


1
Expert's answer
2022-02-28T11:01:51-0500

The molecular formula of ammonium sulphate is (NH4)2SO4

one molecule of ammonium sulphate has 2 N atom.

so,one mole of ammonium sulphate is 2 mole of N atom and ha 2× 6.023×1023 N atom.

Total number of N atom= 6.0×1025 atom

Total number of moles= \frac{6.0×1025 }{12×1023}= 50 mole

so mole of ammonium sulphate is 25 mole.